# Burgers equation from non-thermal stationary states in nearly-integrable gases

### v1

@@ Abstract
When a gas of particles interacts with much a larger reservoir the dynamics of density on large scales is typically governed by diffusion. We study this paradigmatic problem for weakly coupled integrable systems and show that this picture is altered when transport is investigated on top of long-lived non-thermal states. Remarkably, for states non-invariant under parity we find Burgers equation arising in the hydrodynamic limit. We explicitly compute the diffusion constant and nonlinear Euler-scale coupling of the Burgers equation using a variant of the Chapman-Enskog theory. We find excellent agreement between our theory and numerical simulations of a simplified model of stochastic two-body collisions, which we call velocity swap models. Our conclusions are based only on Galilean invariance, existence of a small system-bath coupling parameter and a small momentum exchange between the system and the bath.
@@ I Introduction
The non-equilibrium dynamics of low dimensional quantum many-body systems attracts a lot of attention due to recent advances in experimental, numerical and analytical understanding of fundamental phenomena such as thermalisation and transport [1, 2, 3, 4, 5, 6, 7, 8, 9, 10, 11, 12, 13, 14]. A special role in this area is played by the family of integrable models, for which there has been recently a lot of progress in understanding their equilibration and emergent hydrodynamics, dubbed generalized hydrodynamics (GHD) [15, 16, 17]. The developed techniques cover also situations in which the integrability is weakly broken [18, 19, 20, 21, 22, 23, 24, 25, 26, 27, 28, 29, 30].
In such case the resulting theory is similar to a kinetic theory of dilute gases [31]. Namely, the state of the system is described by a local distribution of quasiparticles which propagate with state-dependent effective velocities and occasionally scatter. Additionally, due to presence of integrable interactions there is a diffusion effect [32, 33, 34, 35, 36]. The scattering between the quasiparticles opens a possibility for a thermalisation channel which otherwise is absent. Another thermalisation mechanism is through introducing an external potential [37, 38, 39, 40] but to simplify the problem we do not consider it here.
@@
The observation that integrability breaking modifies the dynamics of quasiparticles similarly as weak interactions alter the dynamics of free particles can be exploited by bringing tools from the kinetic theory to study hydrodynamics of nearly integrable quantum systems. An example of this is application of the Chapman-Enskog theory to understand the emergence of conventional hydrodynamics along with computation of transport coefficients in such systems [41, 42]. The latter was important to establish an existence of parametrically large space-time regime [43, 44, 45, 46] when the conventional hydrodynamics works before giving place for anomalous one [47, 48, 49, 50]. This kinetic picture can be also lifted from the quasiparticles distribution to their correlations leading to a generalization of the famous Born-Bogoliubov-Green-Kirkwood-Yvon (BBGKY) hierarchy [30]. The generalized BBGKY (gBBGKY) provides a new theoretical framework to understand the integrability-breaking-induced thermalization and was proven useful in description of long-range interacting hard rod gas numerics as well as experiments in 1d dipolar gases [51].
In this work we continue following that program and consider a canonical problem of diffusion of Galilean particles. The mechanism for diffusion is provided by a long-range coupling to a much larger bath, also assumed to be described by a one-dimensional integrable model. As the particles of the smaller system, due to interactions, exchange momentum and energy with the bath, the only conserved quantity is their total number. On a phenomenological level, when the bath state is assumed to be thermal, the continuity equation for the density $\varrho(x,t)$, combined with the gradient expansion of the current, lead to the diffusion equation
@@
with the density dependent diffusion coefficient $\mathcal{D}(\varrho)$. The standard form of the diffusion equation is obtained by linearising the theory with the diffusion coefficient attaining a constant value. In such case the diffusion coefficient can be also computed through the Green-Kubo formula from the velocity autocorrelation function.
The diffusion equation can be also obtained from microscopic considerations based on the kinetic theory and Boltzmann collision operator $\mathcal{I}[\rho_{\mathrm{p}},\rho_{\mathrm{p}}^{(b)}]$ for system-bath interactions. Here $\rho_{\mathrm{p}}(x,\lambda,t)$ is one-body density function ($\lambda$ labels momentum) and $\rho_{\mathrm{p}}^{(b)}$ denotes thermal state of the bath. The density is $\varrho(x,t)=\int{\rm d}\lambda\,\rho_{\mathrm{p}}(x,\lambda,t)$ and under the assumption $\varrho\ll\varrho^{(b)}$ the dynamics of the bath is effectively frozen out. The diffusion coefficient can be computed with the Chapman-Enskog method, discussed later in the text. It is given by an integral equation formulated in terms of the collision operator $\mathcal{I}$.
Our results generalize in various way the diffusion equation (1). Firstly, the resulting diffusion coefficient incorporates in non-perturbative way the integrable interactions which extends the applicability of the standard kinetic theory methods to strongly correlated systems. Secondly, isolated integrable systems do not equilibrate to thermal states due to presence of an extensive number of conservation laws. While breaking the integrability through system-bath interactions ultimately thermalizes the system, if it is sufficiently weak it creates a time span in which the system might be in a prethermal quasi-stationary state [52, 30, 53]. This leads to a new physical situation in which the particles diffuse in a non-thermal environment. Finally, if the non-thermal state is furthermore not parity invariant there are new terms appearing in the diffusion equation. These together ultimately lead to the Burgers equation in the form
@@
where $u(\varrho)$ is advective coefficient, which can be constant or not, depending on particles’ statistics. For classical statistics, $u(\varrho)$ is constant, and equivalent to a simple Galilean boost of the diffusion equation (1). The more interesting situation happens for quantum statistics, where $u(\varrho)$ is non-constant and we recover a genuine Burgers equation. This is summarized in Tab. 1.
@@
symmetric diff. eq. (1) diff. eq. (1)
asymmetric diff. eq. (1) Burgers eq. (2)
Table 1: Hydrodynamic equations for the system for different particle statistics and symmetries of the stationary state.
Equation (2) for the almost homogeneous density profile, $\varrho\approx\varrho_{0}+\delta\varrho$, after changing the reference frame to the one moving with $u(\varrho_{0})$, simplifies to
@@
with time measured in units of $u^{\prime}(\varrho_{0})$ and $\nu=\mathcal{D}(\varrho_{0})/u^{\prime}(\varrho_{0})$. In the literature, it is equation (3) that is usually called Burgers’ equation. Here, for simplicity, we will refer to both equations (2) and (3) as Burgers’ equations.
We exemplify our approach by considering three different scenarios. First, we study a system of two coupled 1d Bose gases which mimics the experimental settings of cold atomic gases consisting of arrays of 1d systems with possibly tunable couplings [51]. It is known that the collision integral for such systems supports athermal stationary states [52]. This provides a concrete physical realisation for testing the ideas presented in this work. Next, we abstract from a concrete realization by considering a generic collision integral in the relaxation time approximation (RTA) with respect to athermal state and with particle number being the only conserved quantity. Within the RTA the Chapman-Enskog equation for the diffusion constant can be solved. This allows us to draw general conclusions for diffusion processes in such circumstances and show how Burgers equations arise for athermal and parity-breaking background states.
Finally, we consider a simple stochastic model of classical or quantum particles colliding with bath particles at constant rate. This serves as a minimal model in which all the phenomena seen in a more sophisticated situations are present but without technical complications. This allows us to pinpoint the sufficient conditions required to see the Burgers’ equations. At the same time this model can be simulated efficiently and we test the effective theory given by (3) against the molecular dynamics. The paper is organized in the following way. In Section II we provide the necessary ingredients of the kinetic theory of nearly integrable system, specifically discuss in the details the GHD-Boltzmann equation and the generalized Gibbs ensemble formulation of the quasi-particle distribution function. In Section III we discuss the collision integrals supporting athermal states. In Section IV we use the Chapman-Enskog method to derive from the GHD-Boltzmann equation the Burgers equation. In Section V, we extract the characteristic relaxation time from the collision integral for two coupled 1d gases which then serves as the input for the relaxation time approximation (RTA). We compute the diffusion coefficient in the latter approximation and link it with thermodynamic properties of the gas also in athermal states. We also show how Burgers equation arises for states without parity invariance. In Section VI we demonstrate that our predictions match the microscopic simulations of the molecular dynamics. We relegated some more technical aspects of our work to the appendices, to which we refer when needed.
@@ II Preliminaries
We consider a system described by a Hamiltonian
$\hat{H}=\hat{H}_{\mathrm{int}}+\hat{H}_{\mathrm{int}}^{(b)}+\alpha\hat{H}_{\mathrm{pert}},$ (4)
where $\hat{H}_{\mathrm{int}}$ and $\hat{H}_{\mathrm{int}}^{(b)}$ are some integrable Hamiltonians modelling system and bath, respectively and $\hat{H}_{\mathrm{pert}}$ is some small integrability-breaking perturbation Hamiltonian (modelling system-bath interaction). The system is assumed to be Galilean-invariant. We consider only the case when $\hat{H}_{\mathrm{pert}}$ does not depend on position and time explicitly. The dynamics of nearly-integrable models, after the initial decoherence phase, is effectively described by a kinetic theory [53, 30, 36, 20]. This kinetic theory is formulated in terms of the space-time dependent distribution function of quasi-particles $\rho_{\mathrm{p}}(x,\lambda,t)$ with quasi-momentum (rapidity) $\lambda$. To understand the basics of quasiparticle description we review first the thermodynamic Bethe ansatz (TBA) and GHD of integrable models.
@@ II.1 Basics of TBA and GHD
The integrable systems are characterized by infinite number of local conserved charges $\hat{Q}_{k}$ fulfilling $H_{\rm int},\hat{Q}_{k}]=0$. Their eigenvalues for a state described by quasiparticle distribution $\rho_{\mathrm{p}}$ are
$Q_{k}=\int{\rm d}\lambda\,h_{k}(\lambda)\rho_{\mathrm{p}}(\lambda).$ (5)
For Galilean invariant models of particles with unit mass, as is assumed here, it is convenient to choose $h_{k}(\lambda)=\lambda^{k}/k!$ which implies that $\hat{Q}_{k}$ are ultra-local charges. Specifically, for total momentum $h_{1}(\lambda)=\lambda$ and for total energy $h_{2}(\lambda)=\lambda^{2}/2$. The quasiparticle picture is based upon the assumption of relaxation to Generalized Gibbs Ensemble (GGE) [54, 55] with the density matrix $\hat{\rho}_{\rm GGE}\propto\exp(-\sum_{j}\beta_{j}\hat{Q}_{j})$, where $\beta_{j}$ are Lagrange multipliers.
We introduce now more ingredients of the TBA in simlar way as presented in [56]. We start with free energy density $\mathsf{f}=\frac{1}{2\pi}\int{\rm d}\lambda F[\epsilon(\lambda)]$, where $\epsilon(\lambda)$ is pseudoenergy and function $F$ depends on particles’ statistics (for fermionic particles $F[\epsilon(\lambda)]=-\log(1+e^{-\epsilon(\lambda)})$). The pseudoenergy is determined from the following equation
@@
with bare pseudoenergy given by $\epsilon_{0}(\lambda)=\sum_{j}\beta_{j}h_{j}(\lambda)$. The set $\{\beta_{j}\}_{j=0,1,\ldots}$ specifies the state. From pseudoenergy we can compute statistical factor $f(\lambda)=({\rm d}^{2}F/{\rm d}\epsilon^{2})/({\rm d}F/{\rm d}\epsilon)$ and occupation function $n={\rm d}F/{\rm d}\epsilon$ defined also as $n=\rho_{\mathrm{p}}/\rho_{\rm t}$, where $\rho_{\rm t}$ is the total density of states. It fulfills the following integral equation
@@
The big Dressing is the result of the action of the operator $\mathbb{F}$, $g^{\rm Dr}=\mathbb{F}^{t}g$ with
@@
Above we introduced backflow function $F(\lambda|\mu)$ [57]. The two dressings are related
@@
For future reference we introduce Dressed momentum $k(\lambda)=(h_{1})^{\rm Dr}(\lambda)$ and energy $\omega(\lambda)=(h_{2})^{\rm Dr}(\lambda)$.
@@
Alternatively, the effective velocity can be written as $v(\lambda)=(h_{1})^{\rm dr}(\lambda)/(h_{0})^{\rm dr}(\lambda)$. For later use it will be convenient to introduce kernels of hydrodynamic matrices: flux Jacobian
$\mathcal{A}=(1-n\mathcal{T})^{-1}v(1-n\mathcal{T})^{-1},$ (12)
along with charge and current susceptibility matrices
$\mathcal{C}$ $=\left(1-n\mathcal{T}\right)^{-1}\rho_{\mathrm{p}}f\left(1-\mathcal{T}n\right)^{-1},$ (13)
$\mathcal{B}$ $=\left(1-n\mathcal{T}\right)^{-1}v\,\rho_{\mathrm{p}}f\left(1-\mathcal{T}n\right)^{-1}.$ (14)
The last ingredient is the diffusion operator, which we denote by $\mathfrak{D}_{\rho}=\mathfrak{D}_{\rho}(\lambda,\mu;x,t)$. It is state-dependent and expressible through objects already introduced. For specific form of $\mathfrak{D}$ and discussion of its properties we refer to Appendix C.1.

### v2

@@ Abstract
When a gas of particles interacts with much a larger reservoir the density dynamics on large scales is typically governed by diffusion. We study this paradigmatic problem for weakly coupled integrable systems and show that this picture gets altered, when transport is investigated on top of long-lived non-thermal states. Remarkably, for states non-invariant under parity we find Burgers equation arising in the hydrodynamic limit. We explicitly compute the diffusion constant and nonlinear advective coefficient of the Burgers equation using a variant of the Chapman-Enskog theory. We find an excellent agreement between our theory and numerical simulations of a simplified model of stochastic two-body collisions. Our conclusions are based only on Galilean invariance, existence of a small system-bath coupling parameter and a small momentum exchange between the system and the bath particles during two-body scattering.
@@ I Introduction
The non-equilibrium dynamics of low-dimensional quantum many-body systems attracts a lot of attention due to recent advances in experimental, numerical and analytical understanding of fundamental phenomena such as thermalisation and transport [1, 2, 3, 4, 5, 6, 7, 8, 9, 3, 10, 11, 12, 13, 14, 15, 16, 17, 18, 19, 20]. In one dimension, one often encounters various instances of anomalous, or non-conventional hydrodynamics. This happens for instance in momentum-conserving fluids or anharmonic chains [21, 22, 23, 24], Luttinger liquids [25, 26] and certain integrable spin chains [27, 28, 29, 30]. Non-conventional transport occurs more broadly in integrable systems, due to the presence of infinite number of conservation laws. In the recent years, a special role in the field of 1D transport is indeed played by the family of such models [31, 32], for which there has been recently a lot of progress in understanding their equilibration and emergent hydrodynamics, dubbed generalized hydrodynamics (GHD) [33, 34, 35]. The developed techniques cover also situations in which the integrability is weakly broken [36, 37, 38, 39, 40, 41, 42, 43, 44, 45, 46, 47, 48].
In such case the resulting description is similar to a kinetic theory of dilute gases [49, 50]. Namely, the state of the system is described by a local distribution of quasiparticles which propagate with state-dependent effective velocities and occasionally scatter due to integrability breaking effects. Additionally, because of the presence of interactions in integrable theory there is a quasiparticle diffusion effect [51, 52, 53, 54, 55]. The scattering between the quasiparticles opens a possibility for a thermalisation channel which otherwise is absent. Another thermalisation mechanism is through introducing an external potential [56, 57, 58, 59], but to simplify the problem we do not consider it here.
@@
The observation that integrability breaking modifies the dynamics of quasiparticles similarly as weak interactions alter the dynamics of free particles can be exploited by bringing tools from the kinetic theory. In particular, such approach was used to study hydrodynamics of nearly-integrable quantum systems. An example of this idea is application of the Chapman-Enskog (ChE) theory [49] to understand the emergence of conventional hydrodynamics along with computation of transport coefficients in such systems [60, 61]. The latter was important to establish an existence of parametrically large space-time regime [62, 63, 64, 65, 66, 60], where the conventional hydrodynamics works before giving place for anomalous one [21, 22, 23, 24]. This kinetic picture can be also lifted from the quasiparticles distribution to their many-body correlations leading to a generalization of the famous Born-Bogoliubov-Green-Kirkwood-Yvon (BBGKY) hierarchy [48]. The generalized BBGKY (gBBGKY) provides a new theoretical framework to understand the integrability-breaking-induced thermalization and was proven useful in description of long-range interacting hard rod gas numerics as well as experiments in 1D dipolar gases [67].
In this work we continue following that program and consider a canonical problem of diffusion of Galilean particles. The mechanism for diffusion is provided by a long-range coupling to a much larger bath, also assumed to be described by a one-dimensional integrable model. As shown in Fig. 1, one can naturally think about such scenario in the geometry of mixed systems, cf. Fig. 1(a) or in coupled 1D tubes, cf. Fig 1(b). The particles of the smaller system, due to long-range interactions, exchange momentum and energy with the bath and therefore the only conserved quantity associated with them is their total number. On a phenomenological level, when the bath state is assumed to be thermal, the continuity equation for the particle density $\varrho(x,t)$, combined with the gradient expansion of the current, lead to the diffusion equation
@@
with the density-dependent diffusion coefficient $\mathcal{D}(\varrho)$. The standard form of the diffusion equation is obtained by linearising the theory with the diffusion coefficient attaining a constant value. In such case the diffusion coefficient can be also computed through the Green-Kubo formula from the velocity autocorrelation function.
The diffusion equation can be also obtained from microscopic considerations based on the kinetic theory and Boltzmann collision operator $\mathcal{I}[\rho_{\mathrm{p}},\rho_{\mathrm{p}}^{(b)}]$ for system-bath interactions. Here $\rho_{\mathrm{p}}(x,\lambda,t)$ is one-body density function ($\lambda$ labels quasimomentum) and $\rho_{\mathrm{p}}^{(b)}$ denotes (thermal) state of the bath. The density is $\varrho(x,t)=\int{\rm d}\lambda\,\rho_{\mathrm{p}}(x,\lambda,t)$ and under the assumption $\varrho\ll\varrho^{(b)}$ the dynamics of the bath is effectively frozen out. The diffusion coefficient can be computed with the ChE method, discussed later in the text, in Sec. IV. It is given by an integral equation formulated in terms of the collision operator $\mathcal{I}$.
Our results generalize in various way the diffusion equation (1). Firstly, the resulting diffusion coefficient incorporates in non-perturbative way the interactions of integrable theory, which extends the applicability of the standard kinetic theory methods to strongly correlated systems. Secondly, isolated integrable systems do not equilibrate to thermal states due to presence of an extensive number of conservation laws. While breaking the integrability through system-bath interactions ultimately thermalizes the system, if it is sufficiently weak it creates a time span in which the system might be in a prethermal quasi-stationary state [68, 48, 12, 45, 46]. This leads to a new physical situation in which the particles diffuse in a non-thermal environment. Finally, if the non-thermal state is furthermore not parity invariant there are new terms appearing in the diffusion equation. These together ultimately lead to the celebrated Burgers equation (studied recently in many contexts [69, 70, 71]) in the form
@@
where $u(\varrho)$ is advective coefficient, which can be constant or not, depending on particles’ statistics. For classical statistics, we find that $u(\varrho)$ is constant, and equivalent to a simple Galilean boost of the diffusion equation (1). The more interesting situation happens for quantum statistics, where $u(\varrho)$ is non-constant and we recover a genuine Burgers equation. This is summarized in Tab. 1.
@@
parity-symmetric diff. eq. (1) diff. eq. (1)
parity-asymmetric diff. eq. (1) Burgers eq. (2)
Table 1: Hydrodynamic equations for the system for different particle statistics and properties of the stationary state.
Equation (2) for an almost homogeneous density profile, $\varrho\approx\varrho_{0}+\delta\varrho$, after changing the reference frame to the one moving with $u(\varrho_{0})$, simplifies to
@@
with time rescaled by $u^{\prime}(\varrho_{0})$ and $\nu=\mathcal{D}(\varrho_{0})/u^{\prime}(\varrho_{0})$. In the literature, it is equation (3) that is usually called Burgers equation. Here, for simplicity, we will refer to both equations (2) and (3) as Burgers equations.
We exemplify our approach by considering three different physical scenarios. First, we study a system of two coupled 1D Bose gases which mimics the experimental settings of cold atomic gases consisting of arrays of 1D systems with possibly tunable couplings [67]. It is known that the collision integral for such systems supports athermal stationary states [68]. This provides a concrete physical realisation for testing the ideas presented in this work. Next, we abstract from a concrete realization by considering a generic collision integral in the relaxation time approximation (RTA) with respect to athermal state and with particle number being the only conserved quantity. Within the RTA the ChE equation for the diffusion constant can be solved. This allows us to draw general conclusions for diffusion processes in such circumstances and show how Burgers equations arise for athermal and parity-breaking background states. Finally, we consider a simple stochastic model of classical or fermionic particles colliding with bath particles at constant rate. This serves as a minimal model in which all the phenomena seen in a more sophisticated situations are present, but without technical complications. This allows us to pinpoint the sufficient conditions required to see the Burgers equation. At the same time, this model can be simulated efficiently and we successfully test the effective theory given by (2) against the molecular dynamics.
The paper is organized in the following way. In Sec. II we provide the necessary ingredients of the kinetic theory of nearly-integrable system, specifically we discuss in the details the Thermodynamic Bethe Ansatz (TBA), the GHD equation supplemented with the Boltzmann term and the Generalized Gibbs Ensemble (GGE) formulation of the quasiparticle distribution function. In Section III we discuss the collision integrals supporting athermal states. In Sec. IV we use the ChE method to derive from the GHD-Boltzmann equation the Burgers equation. In Sec. V, we extract the characteristic relaxation time from the collision integral for two coupled 1D gases which then serves as the input for the RTA. We compute the diffusion coefficient in the latter approximation and link it with thermodynamic properties of the gas also in athermal states. We also show how Burgers equation arises for states without parity invariance. In Sec. VI we demonstrate that our predictions match the microscopic simulations of the stochastic models. We relegated some more technical aspects of our work to the Appendices, to which we refer when needed.
@@ II Preliminaries
We consider a system described by the following Hamiltonian
$H=H_{\mathrm{int}}+H_{\mathrm{int}}^{(b)}+\alpha H_{\mathrm{pert}},\qquad\alpha\ll 1,$ (4)
where $H_{\mathrm{int}}$ and $H_{\mathrm{int}}^{(b)}$ are some integrable Hamiltonians modelling system and bath, respectively and $\alpha H_{\mathrm{pert}}$ is some small integrability-breaking perturbation Hamiltonian (modelling system-bath interaction). The system is assumed to be Galilean-invariant. We consider only the case when $H_{\mathrm{pert}}$ does not depend on position and time explicitly. The dynamics of nearly-integrable models, after the initial decoherence phase, is effectively described by kinetic theory [12, 48, 55, 38]. This kinetic theory is formulated in terms of the space-time dependent distribution function of quasiparticles $\rho_{\mathrm{p}}(x,\lambda,t)$ with quasimomentum (rapidity) $\lambda$. To understand the basics of quasiparticle description we review first the TBA and GHD of integrable models.
@@ II.1 Basics of TBA and GHD
Integrable systems are characterized by infinite number of local conserved charges $Q_{k}$ fulfilling $H_{\rm int},Q_{k}]=0$. Their expectation values for a state described by quasiparticle distribution $\rho_{\mathrm{p}}$ are
$\langle Q_{k}\rangle=\int{\rm d}\lambda\,h_{k}(\lambda)\rho_{\mathrm{p}}(\lambda).$ (5)
For Galilean invariant models of particles with unit mass, as is assumed here, it is convenient to choose $h_{k}(\lambda)=\lambda^{k}/k!$ which implies that $Q_{k}$ are ultra-local charges. Specifically, for total momentum $h_{1}(\lambda)=\lambda$ and for total energy $h_{2}(\lambda)=\lambda^{2}/2$. The quasiparticle picture is based upon the assumption of relaxation to GGE [72, 73] with the density matrix $\rho_{\rm GGE}\propto\exp(-\sum_{j}\beta_{j}Q_{j})$, where $\beta_{j}$ are Lagrange multipliers.
We introduce now more ingredients of the TBA in simlar way as presented in [74]. We start with free energy density $\mathsf{f}=\frac{1}{2\pi}\int{\rm d}\lambda F[\epsilon(\lambda)]$, where $\epsilon(\lambda)$ is so-called pseudoenergy and function $F$ depends on particles’ statistics (for fermionic particles $F[\epsilon(\lambda)]=-\log(1+e^{-\epsilon(\lambda)})$). The pseudoenergy is determined from the following equation
@@
with bare pseudoenergy given by $\epsilon_{0}(\lambda)=\sum_{j}\beta_{j}h_{j}(\lambda)$. Moreover, by $\mathcal{T}(\lambda-\lambda^{\prime})$ we denoted model-dependent scattering kernel. The set $\{\beta_{j}\}_{j=0,1,\ldots}$ specifies the state. From pseudoenergy we can compute statistical factor $f(\lambda)=({\rm d}^{2}F/{\rm d}\epsilon^{2})/({\rm d}F/{\rm d}\epsilon)$ and occupation function $n={\rm d}F/{\rm d}\epsilon$ defined also as $n=\rho_{\mathrm{p}}/\rho_{\rm t}$, where $\rho_{\rm t}$ is the total density of states. It fulfills the following integral equation
@@
The big Dressing is the result of the action of the operator $\mathbb{F}$, $g^{\rm Dr}=\mathbb{F}^{t}g$ ($t$ denotes transposition) with
@@
Above we introduced backflow function $F(\lambda|\mu)$ [32], appearing naturally in studies of excitations in integrable systems. The two dressings are related
@@
For future reference we introduce Dressed momentum $k(\lambda)=h_{1}^{\rm Dr}(\lambda)$ and energy $\omega(\lambda)=h_{2}^{\rm Dr}(\lambda)$. We also mention useful relation that $k^{\prime}(\lambda)=2\pi\rho_{\rm t}(\lambda)$.
@@
Alternatively, the effective velocity can be written as $v(\lambda)=h_{1}^{\rm dr}(\lambda)/h_{0}^{\rm dr}(\lambda)$. For later use it will be convenient to introduce kernels of hydrodynamic matrices: flux Jacobian
$\mathcal{A}=(\mathbf{1}-n\mathcal{T})^{-1}v(\mathbf{1}-n\mathcal{T}),$ (12)
along with charge and current susceptibility matrix kernels
$\mathcal{C}$ $=\left(\mathbf{1}-n\mathcal{T}\right)^{-1}\rho_{\mathrm{p}}f\left(\mathbf{1}-\mathcal{T}n\right)^{-1},$ (13)
$\mathcal{B}$ $=\left(\mathbf{1}-n\mathcal{T}\right)^{-1}v\,\rho_{\mathrm{p}}f\left(\mathbf{1}-\mathcal{T}n\right)^{-1}.$ (14)
The last ingredient is the diffusion operator, which we denote by $\mathfrak{D}_{\rho}=\mathfrak{D}_{\rho}(\lambda,\mu;x,t)$. It is state-dependent and expressible through objects already introduced. For specific form of $\mathfrak{D}$ and discussion of its properties we refer to Appendix D.1.